# On the CP Nature of the ‘95 GeV’ Anomalies

### v1

@@ 1 Introduction
Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC. In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs. Benbrik:2022azi ; Benbrik:2022tlg ; Belyaev:2023xnv ; Belyaev:2024lah ; Benbrik:2024ptw that the 2HDM Type-III with a particular Yukawa texture can successfully accommodate both measurements simultaneously with the lightest CP-even Higgs boson of the model, while being consistent with all relevant theoretical and experimental constraints. In fact, also a 2HDM Type-I can afford similar explanations to the excesses Khanna:2024bah . Further recent studies have shown that actually all three aforementioned signatures can be simultaneously explained in the 2HDM plus a real (N2HDM) Biekotter:2022jyr or a complex (S2HDM) Biekotter:2023jld ; Biekotter:2023oen singlet.
@@ 3.4 Event Simulation and Selection
We implemented our simplified model for event simulation in FeynRules-2.3Alloul:2013bka . The signal and the background events are generated using MadGraph5-aMC@NLO-2.6.6Alwall:2011uj ; Alwall:2014hca . To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-2.6.6. For showering and hadronization we used PYTHIA-8.2 Sjostrand:2006za ; Sjostrand:2014zea , and detector effects were simulated using Delphes-3.4.2deFavereau:2013fsa . We use anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV. For $\tau$-tagging we have used medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency. The signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$. We veto events with any $b$-tagged jets.
@@
We obtain the charged and neutral pions from the $\tau$-jet constituents. From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion. Neutral pions deposit their energies in the hadron calorimeter, HCAL. From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion. For the combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau$ ($\tau\to\pi\nu_{\tau}$).

### v2

@@ 1 Introduction
Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Bhattacharya:2023lmu ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn ; Arhrib:2024wjj ; Gao:2024qag have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC. In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs. Benbrik:2022azi ; Benbrik:2022tlg ; Belyaev:2023xnv ; Belyaev:2024lah ; Benbrik:2024ptw that the 2HDM Type-III with a particular Yukawa texture can successfully accommodate both measurements simultaneously with the lightest CP-even Higgs boson of the model, while being consistent with all relevant theoretical and experimental constraints. In fact, also a 2HDM Type-I can afford similar explanations to the excesses Khanna:2024bah . Further recent studies have shown that actually all three aforementioned signatures can be simultaneously explained in the 2HDM plus a real (N2HDM) Biekotter:2022jyr or a complex (S2HDM) Biekotter:2023jld ; Biekotter:2023oen singlet.
@@ 3.4 Event Simulation and Selection
We implemented our simplified model for event simulation in FeynRules-2.3Alloul:2013bka . The signal and the background events are generated using MadGraph5-aMC@NLO-3.5.3Alwall:2011uj ; Alwall:2014hca . To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-3.5.3. For showering and hadronization we used PYTHIA-8.3 Sjostrand:2014zea ; Bierlich:2022pfr , and detector effects were simulated using Delphes-3.5.0deFavereau:2013fsa . We use anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV. For $\tau$-tagging we have used medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency. The signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$. We veto events with any $b$-tagged jets.
@@
We obtain the charged and neutral pions from the $\tau$-jet constituents. From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion. Neutral pions decay promptly into a pair of photons and therefore deposit their energies in the electromagnetic calorimeter, ECAL. From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion. For the combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau$ ($\tau\to\pi\nu_{\tau}$).